# Multilevel Markov Chain Monte Carlo with likelihood scaling

### v2

@@ 4.1.3 Strong convergence of PDE solution
The following theorem shows the strong convergence of the PDE solution in terms of the truncation number of the KL expansion. We will only look at positive transformations $a$ and additionally require a weak upper bound on $a$. This theorem in particular motivates our choice of workaround: while the necessary bounds do hold for transformed distributions where $\varphi>0$ (and that meet the assumptions of this theorem), they do not hold in the limit of $\varphi=0$. However, the validity for any $\varphi>0$ means that we can approximate the distributions, which suffices for many practical applications. In the context of this theorem, $u_{M}$ is the forward solution of the PDE based on a truncated KL expansion after $M$ terms of the Gaussian field $g$. The proof is based on the proof of [8, Theorem 4.2], where similar results are derived for a diffusion equation. The estimates obtained there also hold for the linear elasticity equations 4, as shown in [7, Chapter 11].
@@
$\|u-u_{M}\|_{L^{p}(H_{0}^{1}(\overline{D}),\Omega)}\lesssim M^{-\frac{1}{2}\left(\frac{k}{d}-1\right)+\delta}$
@@
$a^{\min}_{\varphi}(\omega)\geq\dfrac{\varphi}{2}\min\left\{1,e^{-\|g(\omega)\|_{\mathcal{C}^{0}(D)}}\right\}$
@@
$\|u-u_{M}\|_{L^{p}(H_{0}^{1}(\overline{D}),\Omega)}\leq\left\|\dfrac{1}{a^{\min}_{M,\varphi}}\right\|_{L^{q}(\Omega)}\|a_{\varphi}-a_{M,\varphi}\|_{L^{r}(\mathcal{C}^{0}(\overline{D}),\Omega)}\|b\|_{L^{2}}C_{D}\left\|\dfrac{1}{a_{\varphi}^{\min}}\right\|_{L^{s}(\Omega)}$
@@
$\|u-u_{M}\|_{L^{p}(H_{0}^{1}(\overline{D}),\Omega)}\lesssim(R_{M}^{\beta})^{1/2}$
@@
$R_{M}^{\beta}\lesssim\sum_{m>M}|\lambda_{m}|^{1-2s}\lesssim\sum_{m>M}m^{-k/d(1-2s)}\lesssim M^{1-(1-2s)\frac{k}{d}}$
@@ 4.2.1 Main theorem: cost of multilevel estimator
The main result is Theorem 4.10, giving a bound on the $\varepsilon$-cost (i.e. the cost of obtaining a root mean squared error $<\varepsilon$) of the multilevel estimator under a set of assumptions. Referring to the notation introduced in Subsection 3.1.1, in the context of this theorem $\boldsymbol{\Theta}_{\ell}=\{\theta_{\ell}^{n}\}_{n\in\mathbb{N}}\cup\{\Theta_{\ell-1}^{n}\}_{n\in\mathbb{N}}$ for $\ell\geq 1$ and $\boldsymbol{\Theta}_{0}=\{\theta_{0}^{n}\}_{n\in\mathbb{N}}$. $\nu^{\ell,\ell-1}$ is the joint distribution of $\theta_{\ell}$ and $\Theta_{\ell-1}$ for $\ell\geq 1$, where the marginals of $\theta_{\ell}$ and $\Theta_{\ell-1}$ are $\nu^{\ell}$ and $\nu^{\ell-1}$, respectively. $\rho$ is the true distribution of $Q(\vartheta)$, and $\mathbb{E}_{\bullet}$ and $\mathbb{V}_{\bullet}$ denote the expected value and variance of a quantity under the distribution in their subscript, respectively. $\mathcal{C}_{\ell}$ denotes the cost of evaluating the forward model at level $\ell$. Finally, for consistency we let $Y_{0}=Q_{0}$, $\nu^{0,-1}=\nu^{0}$ and $R_{-1}=1$. For more details and a general proof of this theorem, we refer to [15].
@@
    $|\mathbb{E}_{\nu^{\ell}}[Q_{\ell}]-\mathbb{E}_{\rho}[Q]|\lesssim(R_{\ell}^{-\alpha}+M_{\ell}^{-\alpha^{\prime}}),$
@@
    $\mathbb{V}_{\nu^{\ell,\ell-1}}[Y_{\ell}]\lesssim(R_{\ell-1}^{-\beta}+M_{\ell-1}^{-\beta^{\prime}}),$
@@
    $\mathbb{V}_{\boldsymbol{\Theta}_{\ell}}[\widehat{Y}_{\ell,N_{\ell}}^{MC}]+(\mathbb{E}_{\boldsymbol{\Theta}_{\ell}}[\widehat{Y}_{\ell,N_{\ell}}^{MC}]-\mathbb{E}_{\nu^{\ell,\ell-1}}[\widehat{Y}_{\ell,N_{\ell}}^{MC}])^{2}\lesssim N_{\ell}^{-1}\mathbb{V}_{\nu^{\ell,\ell-1}}[Y_{\ell}]$,
@@
    $\mathcal{C}_{\ell}\lesssim R_{\ell}^{\gamma}$,
@@
$\mathcal{C}_{\varepsilon}(\widehat{Q}_{L,\{N_{\ell}\}}^{ML})\lesssim\left\{\begin{array}[]{ll}\varepsilon^{-2}|\log\varepsilon|&\text{if }\ \beta>\gamma,\\ \varepsilon^{-2}|\log\varepsilon|^{3}&\text{if }\ \beta=\gamma,\\ \varepsilon^{-2-(\gamma-\beta)/\alpha}|\log\varepsilon|&\text{if }\ \beta<\gamma.\end{array}\right.$
@@ 4.2.2 Proof of validity of assumptions
    $|\mathcal{F}(v)|\leq C_{\mathcal{F}}\|v\|_{H^{1/2-\delta}}\quad\text{for all }\delta>0$
@@
    $|N_{L}^{-1}\sigma_{F}^{-2}-N_{\ell}^{-1}\sigma_{F,\ell}^{-2}|\lesssim\max\left(M_{\ell}^{-\frac{1}{2}(\frac{k}{d}-1)+\delta},R_{\ell}^{-1/d+\delta}\right)\quad\text{for all }\delta>0$
@@
    The sequence of level-dependent observation functions $\{W_{\ell}:\mathbb{R}^{N_{L}}\to\mathbb{R}^{N_{\ell}}\}_{\ell=0}^{\infty}$ satisfies
    $\|W_{\ell}(u^{\text{obs}})-W^{S}_{\ell}(u^{\text{obs}})\|\lesssim\max\left(M_{\ell}^{-\frac{1}{2}(\frac{k}{d}-1)+\delta},R_{\ell}^{-1/d+\delta}\right)\quad\text{for all }\delta>0$
Assumption A1 is the same as in the original convergence analysis [15]. Assumption A2 now includes the number of data points $N_{\ell}$ used on each level $\ell$, and holds in the same cases as the original level-independent version of this assumption (where $N_{\ell}=N_{L}$ for all $\ell$). Assumption A3 allows for treatments of observations on coarse levels other than simply selecting only those coinciding with nodes on the coarser level, as long as these more general weight functions eventually converge to the identity function as $\ell\to\infty$. Using these three assumptions we can first prove the validity of assumption M1.
@@
$|\mathbb{E}_{\nu^{\ell}}[Q_{\ell}]-\mathbb{E}_{\rho}[Q]|$ $\lesssim\left(R_{\ell}^{-1/d+\delta}+M_{\ell}^{-\frac{1}{2}(\frac{k}{d}-1)+\delta}\right)+d_{\text{Hell}}(\rho,\rho^{\ell})$
@@
$d_{\text{Hell}}^{2}(\rho,\rho^{\ell})\lesssim\mathbb{E}_{\rho_{0}}\left[\left|\dfrac{\|u^{\text{obs}}-F\|^{2}}{N_{L}\sigma_{F}^{2}}-\dfrac{\|W_{\ell}(u^{\text{obs}})-F_{\ell}\|^{2}}{N_{\ell}\sigma_{F,\ell}^{2}}\right|^{2}\right]$
@@
$\hskip 56.9055pt+\|W_{\ell}(u^{\text{obs}})-F_{\ell}\|^{2}\left|N_{L}^{-1}\sigma_{F}^{-2}-N_{\ell}^{-1}\sigma_{F,\ell}^{-2}\right|$
@@
$\mathbb{E}_{\rho_{0}}\left[\|W_{\ell}^{S}(F)-F_{\ell}\|^{q}\right]^{1/q}\lesssim\left(R_{\ell}^{-1/d+\delta}+M_{\ell}^{-\frac{1}{2}(\frac{k}{d}-1)+\delta}\right)$
@@
$\mathbb{E}_{\rho_{0}}\left[\left|\dfrac{\|u^{\text{obs}}-F\|^{2}}{N_{L}\sigma_{F}^{2}}-\dfrac{\|W_{\ell}(u^{\text{obs}})-F_{\ell}\|^{2}}{N_{\ell}\sigma_{F,\ell}^{2}}\right|^{2}\right]\lesssim\left(R_{\ell}^{-1/d+\delta}+M_{\ell}^{-\frac{1}{2}(\frac{k}{d}-1)+\delta}\right)^{2}$
@@ 5 Numerical experiments
We consider synthetic displacement observations, generated from an artificial ground truth for the stiffness field constructed on the finest level of the multilevel hierarchy. This field is obtained by sampling from a $\mathcal{N}(0,I_{M_{L}})$ prior distribution, with $M_{L}$ the truncation number of the KL expansion at the finest level. We assume to have a measurement error $\eta$ corresponding to a fidelity of $\sigma_{F}=10^{-8}$ m, which corresponds to an order of magnitude of about one percent on the observed displacement, which typically on the order of $10^{-6}$ m. The samples are first fed through the KL expansion in Equation 6 and then the transformation of Equation 10 with scale and shape parameters $\mu=0.4$ and $\kappa=2.5$, respectively, to obtain the desired properties for the spatially varying Young’s modulus.
@@ 5.1.2 Experiment results
(19) $\mathbb{E}_{\zeta}\left[\left(1-\alpha_{M_{\ell},R_{\ell}}(\theta_{\ell}^{\prime}|\theta_{\ell}^{n})\right)\right]\lesssim\left(R_{\ell}^{-1/d+\delta}+M_{\ell}^{-\frac{1}{2}(\frac{k}{d}-1)+\delta}\right)$
@@ 5.2 Comparison between single- and multilevel MCMC
The left plot of Figure 5 shows the cost in CPU time of generating an individual sample at each level in this setup. The total cost has two main contributions: solving the linear system to obtain $F_{\ell}(\theta_{\ell})$, and evaluating the likelihood of Equation 18. The black line is the reference cost of a single-level method, where both parts are performed at the finest, most expensive level for each sample. In red is the cost for a direct application of the multilevel MCMC algorithm without the level-dependent treatment of data outlined in Subsection 3.2. Instead, the output $F_{\ell}(\theta_{\ell})$ is interpolated onto the finest-level mesh to be directly compared to $u^{\text{obs}}$. In this case, there is a clear reduction in sample cost at the middle levels $\ell=3,4$, because at these levels solving the forward model is the largest computational bottleneck. However, the cost reduction plateaus at the coarsest levels, because here the cost of solving the forward model drops below that of evaluating the likelihood function, which becomes the new bottleneck. This loss of efficiency is countered by the level-dependent data treatment shown in blue, where the lower cost is continued for the coarsest levels as well.
@@ 5.3 Ground truth reconstruction
As a final experiment, we show the posterior estimates of a given stiffness field using the multilevel MCMC algorithm. The ground truth is generated by sampling 300 standard normal realisations for the $\{\xi_{m}\}$ which are fed first through the KL expansion and then the transformation of Equation 10. This ground truth is shown in the left plot of Figure 6. We run the multilevel MCMC routine with the same values of $R_{\ell},\tau_{\ell}$ as in Section 5.2, and this time $M_{\ell}=20\cdot(\ell+1)$. The resulting posterior mean for this field in shown in the right plot of Figure 6.

### v3

@@ 4.1.3 Strong convergence of PDE solution
The following theorem shows the strong convergence of the PDE solution in terms of the truncation number of the KL expansion. We will only look at positive transformations $a$ and additionally require a weak upper bound on $a$. This theorem in particular motivates our choice of workaround: while the necessary bounds do hold for transformed distributions where $\varphi>0$ (and that meet the assumptions of this theorem), they do not hold in the limit of $\varphi=0$. However, the validity for any $\varphi>0$ means that we can approximate the distributions, which suffices for many practical applications. In the context of this theorem, $u_{M}$ is the forward solution of the PDE based on a truncated KL expansion after $M$ terms of the Gaussian field $g$. The proof is based on the proof of [8, Theorem 4.2], where similar results are derived for a diffusion equation. The estimates obtained there also hold for the linear elasticity equations (4), as shown in [7, Chapter 11].
@@
$\|u-u_{M}\|_{L^{p}(H_{0}^{1}(\overline{D}),\Omega)}\lesssim M^{-\frac{1}{2}\left(\frac{k}{d}-1\right)+\delta}.$
@@
$a^{\min}_{\varphi}(\omega)\geq\dfrac{\varphi}{2}\min\left\{1,e^{-\|g(\omega)\|_{\mathcal{C}^{0}(D)}}\right\}.$
@@
$\|u-u_{M}\|_{L^{p}(H_{0}^{1}(\overline{D}),\Omega)}\leq\left\|\dfrac{1}{a^{\min}_{M,\varphi}}\right\|_{L^{q}(\Omega)}\|a_{\varphi}-a_{M,\varphi}\|_{L^{r}(\mathcal{C}^{0}(\overline{D}),\Omega)}\|b\|_{L^{2}}C_{D}\left\|\dfrac{1}{a_{\varphi}^{\min}}\right\|_{L^{s}(\Omega)},$
@@
$\|u-u_{M}\|_{L^{p}(H_{0}^{1}(\overline{D}),\Omega)}\lesssim(R_{M}^{\beta})^{1/2}.$
@@
$R_{M}^{\beta}\lesssim\sum_{m>M}|\lambda_{m}|^{1-2s}\lesssim\sum_{m>M}m^{-k/d(1-2s)}\lesssim M^{1-(1-2s)\frac{k}{d}}.$
@@ 4.2.1 Main theorem: cost of multilevel estimator
The main result is Theorem 4.10, giving a bound on the $\varepsilon$-cost (i.e. the cost of obtaining a root mean squared error $<\varepsilon$) of the multilevel estimator under a set of assumptions. Referring to the notation introduced in Subsection 3.1.1, in the context of this theorem $\boldsymbol{\Theta}_{\ell}=\{\theta_{\ell}^{n}\}_{n\in\mathbb{N}}\cup\{\Theta_{\ell-1}^{n}\}_{n\in\mathbb{N}}$ for $\ell\geq 1$ and $\boldsymbol{\Theta}_{0}=\{\theta_{0}^{n}\}_{n\in\mathbb{N}}$. $\nu^{\ell,\ell-1}$ is the joint distribution of $\theta_{\ell}$ and $\Theta_{\ell-1}$ for $\ell\geq 1$, where the marginals of $\theta_{\ell}$ and $\Theta_{\ell-1}$ are $\nu^{\ell}$ and $\nu^{\ell-1}$, respectively. $\rho$ is the true distribution of $Q(\vartheta)$, and $\mathbb{E}_{\bullet}$ and $\mathbb{V}_{\bullet}$ denote the expected value and variance of a quantity under the distribution in their subscript, respectively. $\mathcal{C}_{\ell}$ denotes the cost of evaluating the forward model at level $\ell$. For any subscript, $C_{\bullet}$ denotes a real-valued number independent of $\ell$. Finally, for consistency we let $Y_{0}=Q_{0}$, $\nu^{0,-1}=\nu^{0}$ and $R_{-1}=1$. For more details and a general proof of this theorem, we refer to [15].
@@
    $|\mathbb{E}_{\nu^{\ell}}[Q_{\ell}]-\mathbb{E}_{\rho}[Q]|\leq C_{\text{M1}}\ (R_{\ell}^{-\alpha}+M_{\ell}^{-\alpha^{\prime}}),$
@@
    $\mathbb{V}_{\nu^{\ell,\ell-1}}[Y_{\ell}]\leq C_{\text{M2}}\ (R_{\ell-1}^{-\beta}+M_{\ell-1}^{-\beta^{\prime}}),$
@@
    $\mathbb{V}_{\boldsymbol{\Theta}_{\ell}}[\widehat{Y}_{\ell,N_{\ell}}^{MC}]+(\mathbb{E}_{\boldsymbol{\Theta}_{\ell}}[\widehat{Y}_{\ell,N_{\ell}}^{MC}]-\mathbb{E}_{\nu^{\ell,\ell-1}}[\widehat{Y}_{\ell,N_{\ell}}^{MC}])^{2}\leq C_{\text{M3}}\ N_{\ell}^{-1}\mathbb{V}_{\nu^{\ell,\ell-1}}[Y_{\ell}]$,
@@
    $\mathcal{C}_{\ell}\leq C_{\text{M4}}\ R_{\ell}^{\gamma}$,
@@
$\mathcal{C}_{\varepsilon}(\widehat{Q}_{L,\{N_{\ell}\}}^{ML})\leq C_{\text{ML}}\left\{\begin{array}[]{ll}\varepsilon^{-2}|\log\varepsilon|&\text{if }\ \beta>\gamma,\\ \varepsilon^{-2}|\log\varepsilon|^{3}&\text{if }\ \beta=\gamma,\\ \varepsilon^{-2-(\gamma-\beta)/\alpha}|\log\varepsilon|&\text{if }\ \beta<\gamma.\end{array}\right.$
@@ 4.2.2 Proof of validity of assumptions
    $|\mathcal{F}(v)|\leq C_{\mathcal{F}}\|v\|_{H^{1/2-\delta}}\quad\text{for all }\delta>0.$
@@
    $|N_{L}^{-1}\sigma_{F}^{-2}-N_{\ell}^{-1}\sigma_{F,\ell}^{-2}|\lesssim\max\left(M_{\ell}^{-\frac{1}{2}(\frac{k}{d}-1)+\delta},R_{\ell}^{-1/d+\delta}\right)\quad\text{for all }\delta>0.$
@@
    There exists $C_{W}\in\mathbb{R}$ such that the sequence of level-dependent observation functions $\{W_{\ell}:\mathbb{R}^{N_{L}}\to\mathbb{R}^{N_{\ell}}\}_{\ell=0}^{L}$ satisfies
    $\|W_{\ell}(u^{\text{obs}})-W^{S}_{\ell}(u^{\text{obs}})\|\leq C_{W}\max\left(M_{\ell}^{-\frac{1}{2}(\frac{k}{d}-1)+\delta},R_{\ell}^{-1/d+\delta}\right)\quad\text{for all }\delta>0.$
Assumption A1 is the same as in the original convergence analysis [15]. Assumption A2 now includes the number of data points $N_{\ell}$ used on each level $\ell$. As $N_{\ell}=N_{L}$ for $\ell\geq L$, it holds in the same cases as the original level-independent version of this assumption (where $N_{\ell}=N_{L}$ for all $\ell$). Assumption A3 allows for treatments of observations on coarse levels other than simply selecting only those coinciding with nodes on the coarser level, as long as these more general weight functions do not deviate too strongly from $W_{\ell}^{S}$. Using these three assumptions we can first prove the validity of assumption M1.
@@
$|\mathbb{E}_{\nu^{\ell}}[Q_{\ell}]-\mathbb{E}_{\rho}[Q]|$ $\leq C_{Q}\left(R_{\ell}^{-1/d+\delta}+M_{\ell}^{-\frac{1}{2}(\frac{k}{d}-1)+\delta}\right)+d_{\text{Hell}}(\rho,\rho^{\ell}).$
@@
$d_{\text{Hell}}^{2}(\rho,\rho^{\ell})\leq\mathbb{E}_{\rho_{0}}\left[\left|\dfrac{\|u^{\text{obs}}-F\|^{2}}{N_{L}\sigma_{F}^{2}}-\dfrac{\|W_{\ell}(u^{\text{obs}})-F_{\ell}\|^{2}}{N_{\ell}\sigma_{F,\ell}^{2}}\right|^{2}\right].$
@@
$\hskip 56.9055pt+\|W_{\ell}(u^{\text{obs}})-F_{\ell}\|^{2}\left|N_{L}^{-1}\sigma_{F}^{-2}-N_{\ell}^{-1}\sigma_{F,\ell}^{-2}\right|.$
@@
$\mathbb{E}_{\rho_{0}}\left[\|W_{\ell}^{S}(F)-F_{\ell}\|^{q}\right]^{1/q}\leq C_{\text{A1}}\left(R_{\ell}^{-1/d+\delta}+M_{\ell}^{-\frac{1}{2}(\frac{k}{d}-1)+\delta}\right).$
@@
$\mathbb{E}_{\rho_{0}}\left[\left|\dfrac{\|u^{\text{obs}}-F\|^{2}}{N_{L}\sigma_{F}^{2}}-\dfrac{\|W_{\ell}(u^{\text{obs}})-F_{\ell}\|^{2}}{N_{\ell}\sigma_{F,\ell}^{2}}\right|^{2}\right]\leq C_{\text{Hell}}\left(R_{\ell}^{-1/d+\delta}+M_{\ell}^{-\frac{1}{2}(\frac{k}{d}-1)+\delta}\right)^{2},$
@@ 5 Numerical experiments
We consider synthetic displacement observations, generated from an artificial ground truth for the stiffness field constructed on the finest level of the multilevel hierarchy. This field is obtained by sampling from a $\mathcal{N}(0,I_{M_{L}})$ prior distribution, with $M_{L}$ the truncation number of the KL expansion at the finest level. We assume to have a measurement error $\eta$ corresponding to a fidelity of $\sigma_{F}=10^{-8}$ m, which corresponds to an order of magnitude of about one percent on the observed displacement, which typically on the order of $10^{-6}$ m. The samples are first fed through the KL expansion in Equation (6) and then the transformation of Equation (10) with scale and shape parameters $\mu=0.4$ and $\kappa=2.5$, respectively, to obtain the desired properties for the spatially varying Young’s modulus.
@@ 5.1.2 Experiment results
(19) $\mathbb{E}_{\zeta}\left[\left(1-\alpha_{M_{\ell},R_{\ell}}(\theta_{\ell}^{\prime}|\theta_{\ell}^{n})\right)\right]\lesssim\left(R_{\ell}^{-1/d+\delta}+M_{\ell}^{-\frac{1}{2}(\frac{k}{d}-1)+\delta}\right),$
@@ 5.2 Comparison between single- and multilevel MCMC
The left plot of Figure 5 shows the cost in CPU time of generating an individual sample at each level in this setup. The total cost has two main contributions: solving the linear system to obtain $F_{\ell}(\theta_{\ell})$, and evaluating the likelihood of Equation (18). The black line is the reference cost of a single-level method, where both parts are performed at the finest, most expensive level for each sample. In red is the cost for a direct application of the multilevel MCMC algorithm without the level-dependent treatment of data outlined in Subsection 3.2. Instead, the output $F_{\ell}(\theta_{\ell})$ is interpolated onto the finest-level mesh to be directly compared to $u^{\text{obs}}$. In this case, there is a clear reduction in sample cost at the middle levels $\ell=3,4$, because at these levels solving the forward model is the largest computational bottleneck. However, the cost reduction plateaus at the coarsest levels, because here the cost of solving the forward model drops below that of evaluating the likelihood function, which becomes the new bottleneck. This loss of efficiency is countered by the level-dependent data treatment shown in blue, where the lower cost is continued for the coarsest levels as well.
@@ 5.3 Ground truth reconstruction
As a final experiment, we show the posterior estimates of a given stiffness field using the multilevel MCMC algorithm. The ground truth is generated by sampling 300 standard normal realisations for the $\{\xi_{m}\}$ which are fed first through the KL expansion and then the transformation of Equation (10). This ground truth is shown in the left plot of Figure 6. We run the multilevel MCMC routine with the same values of $R_{\ell},\tau_{\ell}$ as in Section 5.2, and this time $M_{\ell}=20\cdot(\ell+1)$. The resulting posterior mean for this field in shown in the right plot of Figure 6.